OpenStudy (doughnuttable):

What does it mean to do something like \( \frac{dx}{dx} x\) rather than just \( \frac{d}{dx} x\) ?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle dy/dx }\) = you're differentiating the function \(\color{#000000 }{ \displaystyle y }\), with respect to \(\color{#000000 }{ \displaystyle x }\). \(\color{#000000 }{ \displaystyle \frac{dx}{dx} }\) = you're differentiating x with respect to x. (d/dx of x is 1). So, \(\color{#000000 }{ \displaystyle \frac{dx}{dx}x }\), would likely be determined as: \(\rm «\)the derivative of x with respect to x, times x\(\rm »\).

OpenStudy (solomonzelman):

I haven't ever seen anyone use: \(\color{#000000 }{ \displaystyle \frac{dx }{dx} x }\) I'm sure enough that this is wrong.

OpenStudy (solomonzelman):

Where as \(\color{#000000 }{ \displaystyle \frac{d }{dx} x }\) = derivative of x with respect to x.

OpenStudy (doughnuttable):

Is \(\frac{dy}{dx}\) the same as \(\frac{d}{dx} y\)?

OpenStudy (solomonzelman):

Yes, that is same, I think. Both mean, \(«{\tiny~}\)the derivative of \(y\) with respect to \(x\)\({\tiny~}»\).